OpenStudy (majikdusty):

A 12-m long ladder leans against a vertical building. If the top of the ladder is sliding down the building at .25m/s, how fast is the bottom of the ladder sliding away from the building when the top of the ladder is 6m above the ground?

OpenStudy (anonymous):

looks like a related rate pythagoras problem

OpenStudy (anonymous):



OpenStudy (anonymous):

\[x^2+y^2=12^2\] \[2xx'+2yy'=0\] \[yy'=-xx'\] and you are told \(x'=.25\) so plug the number in, solve for \(y'\)

OpenStudy (majikdusty):

Thank you very much! For some reason i was using .25 ad dy/dt instead of dx/dt! I appreciate the help! the answer i got was 1/(4sqrt(3))

OpenStudy (anonymous):

oh i made a tiny mistake. it should be \(x'=-.25\) because the ladder is sliding down the wall, and so it is decreasing, making the rate of change negative